OpenStudy (anonymous):

What is the 19th term of the arithmetic sequence 14, 22, 30, … ?

OpenStudy (anonymous):

14+18 \times 8 = 14 + 144 = 158

OpenStudy (anonymous):

\[ 14+18 \times 8 = 14 + 144 = 158 \]

OpenStudy (anonymous):

6+8x where x is the term.

OpenStudy (anonymous):

6+8(19) 6+152 158

OpenStudy (anonymous):

The 19th term is therefor 158 :)

OpenStudy (anonymous):

@madscientist101: Can you explain?

OpenStudy (anonymous):

thank you!! totally diff than how my teacher showed me haha your way is easier lol

OpenStudy (anonymous):

*6+8x @madscientist101

OpenStudy (anonymous):

that's what I wrote FFM 6+8x

OpenStudy (anonymous):

\[T _{n}=a+(n-1)d\] a=14 \[T _{2}=a+(2-1)d=22\] a+d=22 d=22-14=8 \[T _{19}=14+(19-1)8=158\]

OpenStudy (anonymous):

Okay, so you go 1 term back and find the 20th term instead of 19th right? @madscientist101

OpenStudy (anonymous):

The difference between each term is 8. It starts with 14 so it must be 14+8(x-1)... because the you don't want the 14 to be counted as the first term.. although it is... 14+8(x-1) = 14+8x-8 =6+8x @Foolformath

OpenStudy (anonymous):

lol nice Mad :)

OpenStudy (anonymous):

lol yup